(c) If by 31st October 1986, Kendwa Island on Indian Ocean had the magnetic variation of 8010’E and a true bearing of 240030’ from station x, the annual magnetic declaration decreased by 15’. Determine
(a) Magnetic bearing of the Island after 36 years.
(b) The magnetic variation of the Island by 31st October 2018.
1
Expert's answer
2020-04-02T11:59:49-0400

Magnetic variation - the angle between true (geographical) north and the magnetic north.

True (magnetic) bearing - the angle measured clockwise from true north to the object.

Annual magnetic declination - the change in degrees per a year.


Therefore, magnetic bearing of the Island after 36 years will be


"M=80^\circ10'+36\cdot(-0^\circ15')=71^\circ10'."


The magnetic variation of the Island by 31st October 2018:


"M=80^\circ10'+(2018-1986)\cdot(-0^\circ15')=72^\circ10'."